# Consecutive Model Editing with Batch alongside HooK Layers

### v1

@@ 3.1.3 Local Editing Scope Identification
Outlier Demonstration
Given the original outputs produced by the model layer weights and the edited outputs generated by hook layer weights, we need to decide when and which part of the original outputs to swap over. The ideal solution is to only switch those parts of outputs whose keys have been updated to the hook layer weights and leave other parts unchanged. To this end, we first detect the output parts that have their keys updated. Suppose $k_{i}\in K_{1},v_{i}\in V_{1}$ is an association that has been updated to $W_{1}$, and $k_{j}\notin K_{1}$ is a key that is not included in the updated associations. We show empirically in section 4.4 that $\parallel{W_{1}k_{i}-W_{0}k_{i}}\parallel\gg\parallel{W_{1}k_{j}-W_{0}k_{j}}\parallel$ holds. This implies that when the hook layer with updated weight $W_{h}$ receives an input $\hat{K}\in R^{m\times n}$ (batch dimension is ignored for simplicity) that contains an edited key $k_{i}\in\hat{K}\cap K_{1},\,k_{i}\in R^{m}$, then we should have $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel\gg\parallel{W_{h}k_{j}-W_{0}k_{j}}\parallel$ for all $k_{j}\in\hat{K}-\hat{K}\cap K_{1}$, which means that $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel$ would be outliers among $\{\parallel{W_{h}k_{x}-W_{0}k_{x}}\parallel:\!\forall k_{x}\in\hat{K}\}$. Hence, detecting outputs of the updated keys can be transferred to detecting the outliers in the L2-norm distribution of inputs. We used the standardization to find the outliers (Fig.1), which applies the standardization technique to L2-norm vectors of inputs and determines outliers via a predefined threshold $\alpha$. Concretely, for the inputs $\hat{K}$, we first compute the L2-norm vector $M^{l}\in R^{n}$:
@@
$M^{l}_{i}=\parallel(O^{l}_{i}-P^{l}_{i})\parallel\ .$ (17)
Then, we standardize $M^{l}$ and select the swap location by comparing it with $\alpha$.
@@
We denote $Z_{x}=\max((M^{l}-\mu)/\sigma)$ as the maximum z-score of an input $\hat{K}$. Since the $Z_{x}$ varies for different instances (section 4.4) and is likely to shift as the consecutive editing steps grow, it is not reasonable to set $\alpha$ as a fixed real number. Therefore, we determine the $\alpha$ dynamically during the editing process:
$\alpha_{s}=\begin{cases}\alpha_{z}&\text{if}\;s=1,\\ \min(\alpha_{c},\alpha_{s-1})&\text{otherwise}\end{cases}\ ,$ (18)
where $s\geq 1$. Specifically, the $\alpha$ is first initialized to a pre-selected value $\alpha_{z}$. At each consecutive editing step $s$, for the batch of inputs in this step, we calculate $Z_{x}$ for each single instance and select the minimum $Z_{x}$ in the batch (i.e., the supremum) as the candidate $\alpha_{c}$. The $\alpha_{s}$ is finally determined to be the minimum between the candidate $\alpha_{c}$ and the previous value $\alpha_{s-1}$. In practice, we set $\alpha_{z}=2.2$.
Table 1: Single round batch editing results with batch size 30.
Method Model ZsRE COUNTERFACT
Reliability Generality Locality Reliability Generality Locality
FT (Meng et al., 2022) GPT2-XL 0.1685 0.1634 0.7155 0.0027 0.0034 0.8518
MEND (Mitchell et al., 2022a) 0.0357 0.03413 0.1993 0.0013 0.0003 0.0422
SERAC (Mitchell et al., 2022b) 0.4608 0.3007 0.1962 0.0010 0.0045 0.8225
MEMIT (Meng et al., 2023) 0.6119 0.4997 0.9751 0.8101 0.2767 0.9580
COMEBA-HK 0.8221 0.6661 0.9940 0.8828 0.4038 0.9766
FT (Meng et al., 2022) GPT-J 0.2257 0.2177 0.9919 0.0037 0.0034 0.9957
MEND (Mitchell et al., 2022a) 0.1050 0.0918 0.1225 0.1728 0.0351 0.0858
SERAC (Mitchell et al., 2022b) 0.2766 0.2633 0.6052 0.0020 0.0028 0.9999
MEMIT (Meng et al., 2023) 0.9340 0.7045 0.9647 0.9957 0.4229 0.9525
COMEBA-HK 0.9759 0.7241 0.9910 0.8794 0.4276 0.9817
@@ 3.2 Multiple-layer Consecutive Batch Editing
Given the designed single-layer editing procedure, there exists a risk that the single-layer hook fails to detect the updated keys. Suppose $k_{i}$ is an updated key; failure to detect $k_{i}$ indicates that the output corresponds to $k_{i}$ at this single layer would be the same as the original output $W_{proj}k_{i}$, which consequently leads to the failure of edition to $k_{i}$. We believe one potential solution is to apply the hook to multiple model layers rather than a single model layer because the latter layer grasps the chance to capture the edited keys missed by proceeding layers. Furthermore, (Zhu et al., 2020) showcased that minimizing the magnitude of parameter change is helpful for improving the robustness of the model. Thus, we expand our work to multiple layers (Fig.2).
We first find the desired object vector $v_{i}$ following a similar procedure in (Meng et al., 2023). However, the optimization is not based on the original model, but the model hung with the validated hook that inherits the most recently updated hook weights from the previous editing step. After $v_{i}$ is found, we start to update the hook weight at each layer.
At each batch editing step, all the hook layers are initialized to temporary hook layers, which substitute the entire original output to output from hook layers. The purpose of doing this is to ensure that the residual regarding the hook layer weights rather than the original model weights are calculated. Then, the residual is distributed evenly to each layer, and the alteration $\Delta^{l}$ to the parameter at each layer is found in a layer-increasing iterative manner with keys, and residuals recomputed at each iteration (Fig.2). The reason for the recomputation of keys and residuals is that the layer-increasing alteration approach will affect the keys and residuals in the latter layer. For each layer, once the hook layer weight is updated, the hook layer is changed from a temporary hook layer to a validated hook layer to facilitate the computation of the keys and residuals in the latter layer. After the whole editing process is completed, the validated hook layers with the ultimately updated weights are hung on the model to shape the final edited model.
@@ 4.1 Experiment Setups
Datasets
We use the ZsRE (Levy et al., 2017) and COUNTERFACT (Meng et al., 2023) datasets with the split provided in EasyEdit [^1] https://github.com/zjunlp/EasyEdit/tree/main for evaluation. ZsRE is a question-answering dataset that uses back-translation to generate equivalent neighborhoods. It is initially used in factual knowledge evaluation and later adopted in model editing by (Mitchell et al., 2022a). COUNTERFACT is a challenging dataset that focuses on counterfactual information that has a low prediction score compared to correct facts. It builds out-of-scope data by replacing the subject entity with a similar description that shares the same predicate. Further details are provided in Appendix.A.
Baselines
@@
For baselines , we adopt several batch-supportive editing methods, including SERAC (Mitchell et al., 2022b), MEND (Mitchell et al., 2022a), MEMIT (Meng et al., 2023), and fine-tuning (FT) technique used in (Meng et al., 2022; Yao et al., 2023), which only fine-tune a specific layer identified by Rome (Meng et al., 2022) instead of all layers to ensure a fair comparison. We choose large autoregressive language models GPT2-XL and GPT-J (6B) as our base models. All baselines are implemented using the EasyEdit (Yao et al., 2023; Zhang et al., 2024; Wang et al., 2023; Cheng et al., 2023; Mao et al., 2023; Zhang et al., 2023) library. Further details of the datasets and implementation are given in the Appendix A.

### v3

@@ 3.1.3 Local Editing Scope Identification
Outlier Detection
Given the original outputs produced by the model layer weights and the edited outputs generated by hook layer weights, we need to decide when and which part of the original outputs to swap over. The ideal solution is only to switch those parts of outputs whose keys have been updated to the hook layer weights and leave other parts unchanged. To this end, we first detect the output parts that have their keys updated. Suppose $k_{i}\in K_{1},v_{i}\in V_{1}$ is an association that has been updated to $W_{1}$, and $k_{j}\notin K_{1}$ is a key that is not included in the updated associations. We show empirically in section 4.4 that $\parallel{W_{1}k_{i}-W_{0}k_{i}}\parallel\gg\parallel{W_{1}k_{j}-W_{0}k_{j}}\parallel$ holds. This implies that when the hook layer with updated weight $W_{h}$ receives an input $\hat{K}\in R^{m\times n}$ (batch dimension is ignored for simplicity) that contains an edited key $k_{i}\in\hat{K}\cap K_{1},\,k_{i}\in R^{m}$, then we should have $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel\gg\parallel{W_{h}k_{j}-W_{0}k_{j}}\parallel$ for all $k_{j}\in\hat{K}-\hat{K}\cap K_{1}$, which means that $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel$ would be outliers among $\{\parallel{W_{h}k_{x}-W_{0}k_{x}}\parallel:\!\forall k_{x}\in\hat{K}\}$. Hence, detecting outputs of the updated keys can be transferred to detecting the outliers in the L2-norm distribution of inputs. We used the standardization to find the outliers (Fig. 1), which applies the standardization technique to L2-norm vectors of inputs and determines outliers via a predefined threshold $\alpha$. Concretely, for the inputs $\hat{K}$, we first compute the L2-norm vector $M^{l}\in R^{n}$:
@@
$M^{l}=\parallel(O^{l}-P^{l})\parallel$ (17)
Note that $\parallel$.$\parallel$ . ∥ here means computing the L2-norm for each vector over the keys’ dimension ($m$). Then, we standardize $M^{l}$ to get the z-score vector $Z^{l}$ and select the swap location by comparing it with $\alpha$. The details of choosing $\alpha$ are discussed in the next paragraph. Specifically, we do:
$h_{i}^{l}=\begin{cases}O_{i}^{l}&\text{if $Z_{i}^{l}\geq\alpha$},\\ P_{i}^{l}&\text{ortherwise}.\end{cases}$ (18)
where $i$ is the index over tokens.
@@
We denote $Z_{i}^{l}=\max((M^{l}-\mu)/\sigma)$ as the maximum z-score entry of an input $\hat{K}$. Since the $Z_{i}^{l}$ varies for different instances (§4.4) and is likely to shift as the consecutive editing steps grow, it is unreasonable to set $\alpha$ as a fixed real number. Therefore, we determine the $\alpha$ dynamically during the editing process:
$\alpha_{s}=\begin{cases}\alpha_{z}&\text{if}\;s=1,\\ \min(\alpha_{c},\alpha_{s-1})&\text{otherwise}\end{cases}$ (19)
where $s\geq 1$. Specifically, the $\alpha$ is first initialized to a pre-selected value $\alpha_{z}$. At each consecutive editing step $s$, for the batch of inputs in this step, we calculate $Z_{i}^{l}$ (the maximal z-score entry) for each single instance and select the minimal $Z_{i}^{l}$ in the batch (i.e., the supremum) as the candidate $\alpha_{c}$. The $\alpha_{s}$ is finally determined to be the minimum between the candidate $\alpha_{c}$ and the previous value $\alpha_{s-1}$. In practice, we set $\alpha_{z}=2.2$.
Method Model ZsRE COUNTERFACT Reliability Generality Locality Average Reliability Generality Locality Average FT-L Meng et al. (2022) GPT2-XL 16.85 16.34 71.55 34.91 0.27 0.34 85.18 28.60 FT-M 17.95 17.32 71.26 35.51 0.36 0.42 82.81 27.86 LoRA Hu et al. (2022) 30.10 29.08 80.54 46.57 5.64 3.46 69.45 26.18 MEND Mitchell et al. (2022a) 2.16 2.11 20.34 8.20 0.13 0.03 4.22 1.46 SERAC Mitchell et al. (2022b) 98.64 48.12 35.68 60.81 17.88 14.55 82.25 38.23 MEMIT Meng et al. (2023) 61.19 49.97 97.51 69.56 81.01 27.67 95.80 68.16 CoachHooK 82.21 66.61 99.40 82.74 88.28 40.38 97.66 75.44 FT-L Meng et al. (2022) GPT-J 22.57 21.77 99.19 47.84 0.37 0.34 99.57 33.43 FT-M 99.96 80.31 43.35 74.54 99.99 35.29 17.04 50.77 LoRA Hu et al. (2022) 99.97 83.20 17.64 66.93 99.87 53.10 2.50 51.82 SERAC Mitchell et al. (2022b) 87.46 63.64 77.35 76.15 16.67 15.93 99.99 44.20 MEMIT Meng et al. (2023) 93.40 70.45 96.47 86.77 99.57 42.29 95.25 79.04 CoachHooK 97.59 72.41 99.10 89.70 87.94 42.76 98.17 76.29
Table 1: Single round batch editing results. The best two average scores are highlighted. All metrics are desired higher values.
Figure 2: Multiple layer update with hook layer (Attention module and the first layer of FFN are omitted). The value vector $v_{i}$ is first computed at the last editing layer, and then we iteratively insert a fraction of the residual to each editing layer (I, II, III) using Eq. 13. Since changing one layer would affect the activations of downstream layers, recollection of the activations is conducted after each iteration. At the beginning, temporary hook layers are initialized to all editing layers. Once the hook layer weight is updated, it is replaced by the validated hook layer (1, 2, 3).
@@ 3.2 Multiple-layer Consecutive Batch Editing
Given the designed single-layer editing procedure, there exists a risk that the single-layer hook fails to detect the updated keys. Suppose $k_{i}$ is an updated key; failure to detect $k_{i}$ indicates that the output corresponds to $k_{i}$ at this single layer would be the same as the original output $W_{proj}k_{i}$, which consequently leads to the failure update for $k_{i}$. To tackle this issue, one potential solution is to apply the hook to multiple model layers rather than a single model layer because the latter layer grasps the chance to capture the edited keys missed by proceeding layers. Furthermore, Zhu et al. (2020) showcased that minimizing the magnitude of parameter change is helpful for improving the robustness of the model. Thus, we expand our work to multiple layers (Fig. 2).
We first find the desired object vector $v_{i}$ following a similar procedure in Meng et al. (2023). However, the optimization is not based on the original model, but the model hung with the validated hook that inherits the most recently updated hook weights from the previous editing step. After $v_{i}$ is found, the hook weight is updated at each layer.
At each batch editing step, all the hook layers are initialized to temporary hook layers, which substitute the entire original output to output from hook layers. The purpose of doing this is to ensure that the residual regarding the hook layer weights rather than the original model weights are calculated. Then, the residual is distributed evenly to each layer, and the alteration $\Delta^{l}$ to the parameter at each layer is found in a layer-increasing iterative manner with keys and residuals recomputed at each iteration (Fig. 2). The reason for the recomputation of keys and residuals is that the layer-increasing alteration approach will affect the keys and residuals in the latter layer. For each layer, once the hook layer weight is updated, the hook layer is changed from a temporary hook layer to a validated hook layer to facilitate the computation of the keys and residuals in the latter layer. After the whole editing process is completed, the validated hook layers with the ultimately updated weights are hung on the model to shape the final edited model.
@@ 4.1 Experiment Setups
Datasets & Evaluation Metrics
We use the ZsRE Levy et al. (2017) and COUNTERFACT Meng et al. (2023) datasets with the split provided in EasyEdit [^2] https://github.com/zjunlp/EasyEdit/tree/main for evaluation. We employ three popular editing evaluation metrics defined in Yao et al. (2023); Huang et al. (2023); Cao et al. (2021), i.e., Reliability, Generality, and Locality, as well as the average scores [^3] Most of the application scenarios of model editing require good performance in all three metrics. over the three metrics. Further details are provided in Appendix A.
Baselines & Implementation Details
@@
For baselines, we adopt several batch-supportive editing methods, including LoRA Hu et al. (2022), SERAC Mitchell et al. (2022b), MEND Mitchell et al. (2022a), MEMIT Meng et al. (2023) and fine-tuning with specific layer (FT-L) technique used in Meng et al. (2022); Yao et al. (2023), which only fine-tune a specific layer identified by Rome Meng et al. (2022) instead of all layers to ensure a fair comparison. We also include a small variation of FT-L called FT-M and a sequential supportive editing method GRACE Hartvigsen et al. (2022). We choose large autoregressive language models GPT2-XL and GPT-J (6B) as our base models.
For all consecutive editing experiments, the evaluation is conducted after the full set of consecutive steps are finished. For example, in Fig. 4, we conduct experiments for sample sizes 200, 400, 600, 800, and 1000, so the evaluation is triggered right after the first 200, 400, 600, etc, samples are edited to the model. Unless specified, the batch size [^4] Since GRACE Hartvigsen et al. (2022) does not support batch editing, we set the batch size to 1 for GRACE. for consecutive editing is selected to be 10. Further details of the baselines and implementation are given in the Appendix B.
Method Model ZsRE COUNTERFACT Reliability Generality Locality Average Reliability Generality Locality Average FT-L Meng et al. (2022) GPT2-XL 3.79 2.48 6.60 4.29 1.00 1.00 6.00 2.67 FT-M 8.92 8.41 6.22 7.85 4.00 3.50 5.50 4.33 LoRA Hu et al. (2022) 0.96 1.29 0.03 0.76 0.50 0.02 0.50 0.34 MEND Mitchell et al. (2022a) 20.95 18.29 93.69 47.01 0.01 0.00 0.08 0.03 SERAC Mitchell et al. (2022b) 100 36.03 35.95 57.33 15.41 12.96 81.00 36.46 GRACE Hartvigsen et al. (2022) 100 0.04 100 66.68 100 0.40 100 66.80 MEMIT Meng et al. (2023) 34.88 32.96 70.74 46.19 56.00 37.00 31.00 41.33 CoachHooK 66.91 56.11 97.23 73.42 86.00 38.00 59.00 61.00 FT-L Meng et al. (2022) GPT-J 23.53 21.70 55.27 33.5 2.00 2.00 72.00 25.33 FT-M 64.33 55.63 17.59 45.85 25.50 5.00 2.00 10.83 LoRA Hu et al. (2022) 1.43 1.39 0.02 0.95 0.50 0.50 0.10 0.37 SERAC Mitchell et al. (2022b) 86.91 55.36 79.07 73.78 18.49 14.56 98.89 43.98 GRACE Hartvigsen et al. (2022) 100 0.04 100 66.68 100 0.50 100 66.83 MEMIT Meng et al. (2023) 63.36 48.90 74.80 62.35 75.00 45.00 42.00 54.00 CoachHooK 79.89 61.29 96.52 79.23 95.00 41.00 80.00 72.00
Table 2: Consecutive batch editing results.
4.2 Evaluation on Single-round Batch Editing
We first test the effectiveness of our method under basic single-round batch editing settings with batch size 30, i.e., the model is rolled back to the initial state after each batch editing. Both MEMIT and CoachHooK need to set the parameter $\lambda$, the balance factor between pre-trained and newly updated associations. According to Meng et al. (2023), higher $\lambda$ helps preserve the original model behavior (locality) but could harm reliability and generality, and the best overall performance is found at around $\lambda=10^{4}$. However, with the intent to verify whether our method comprehensively improves the editing, that is, could accept lower $\lambda$ to assign higher weight for new associations while not sacrificing the locality, we deliberately set $\lambda=5\times 10^{3}$ for CoachHooK and keep it as the optimized value for MEMIT, which are $2\times 10^{4}$ and $1.5\times 10^{4}$ for GPT2-XL and GPT-J respectively.
The evaluation results are shown in Table 1. For GPT2-XL, our method has the best result in almost every metric. Specifically, despite the relatively low $\lambda$, our method overwhelms other baselines in generality metrics while maintaining a better locality. This indicates that lowering $\lambda$ or, in other words, increasing the weight of the new associations does not sacrifice the locality in CoachHooK. The improvement in GPT-J is less compared with that in GPT2-XL. However, our method still has the best average score for the ZsRE dataset and a comparable average score with the best in the COUNTERFACT dataset.
4.3 Evaluation on Consecutive Batch Editing
We evaluate our method’s capability on 1k samples from both datasets for consecutive batch editing, i.e., there is no roll-back. The evaluation is conducted after the end of the whole consecutive batch editing process. We set $\lambda$ to $15,000$ as the scenario now is consecutive batch editing.
Results in Table 2 show that most of the methods suffer from a great performance drop contrasted to editing in a single round. Although our method’s performance experiences a decrease as well, it surpasses other methods in 100 consecutive steps with an even larger improvement margin for almost all the metrics compared to the single-round batch editing. This demonstrates that our method does not depend on simple regurgitation of the editing samples nor rely heavily on the trade-offs of lowering the balancing factor $\lambda$ to increase the reliability and locality. An interesting point is that the GRACE performs perfectly in reliability and locality but poorly in generality. As expected, GRACE is superior in reliability since it maintains a codebook to memorize the instances of editing that are encountered. However, its inferiority in generality indicates that it suffers from the problem of regurgitation. (One may argue that increasing the deferral radius could help improve the generality of GRACE, we therefore provide further experiments in Appendix C.)
We extend the data scale of the consecutive batch editing experiment to 10k (1k consecutive steps) to explore the limit of our method. Results can be found in Fig. 6. Surprisingly, the locality experiences a great fall from 100 to 200 steps but remains steady from 200 to 1k editing steps, which proves that the hook layer stably obstructs the out-scope samples. Reliability and generality consistently fall as the consecutive steps grow, indicating that there is still room for improvement in this field.
4.4 Validation of Local Editing Scope
Figure 3: Difference between the z-score entry to the updated key $Z^{l}_{key}$ and average of $Z^{l}$. The x-axis represents the sample index.
Given an updated hook layer with the weight $W_{h}$, the original model weight $W_{0}$, an updated key $k_{i}$, and an out-of-scope key $k_{j}$, we conduct experiments to verify whether $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel\gg\parallel{W_{h}k_{j}-W_{0}k_{j}}\parallel$ holds. We select 100 samples from the COUNTERFACT dataset to edit GPT2-XL using CoachHooK, then apply the edited model to these 100 samples and record the z-score entries of the L2-norm of the difference vector between update keys’ response from the last hook layer and original model layer, namely, z-score entries of $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel$. Both reliability and generality prompts are included for comprehensiveness.
The result is shown in Fig. 3. Almost all the z-scores of the responses from updated keys exhibit a great margin from the mean value, with the lowest around 1.5 in reliability samples and 2 in generality samples. The discriminative z-score demonstrates that the identification technique (section 3.1.3) can effectively filter editing-irrelevant instances and accept editing-relevant instances, which validates the local editing scope.
4.5 Detailed Analysis and Discussions
Figure 4: Ablation study.
Ablation Study of Update mechanism and Hook layers
The effectiveness of the derived consecutive updating mechanism and the hook layers are discussed in this part. We run three cases using GPT2-XL, namely, MEMIT (no consecutive updating mechanism, no hook layers), CoachHooK without hook (CoachHooK w/o HK), and CoachHooK for consecutive batch editing on 1k samples from both ZsRE and COUNTERFACT datasets.
The results are demonstrated in Fig. 4. In almost all metrics of the two datasets except the generality of COUNTERFACT, the CoachHooK w/o the hook performs better than the vanilla MEMIT, and the margin tends to increase as the consecutive steps ascend. This certifies the effectiveness of our derived consecutive updating mechanism in consecutive batch editing scenarios. For the ZsRE dataset, the method with hook layers considerably outperforms the one without hook in the locality without sacrificing reliability and generality. This verifies that the hook layer can efficiently and accurately block the out-scope instances from the input without fraudulently missing in-scope instances. For the COUNTERFACT dataset, the reliability of CoachHooK is consistently higher than the other two, and the generality surpasses that of MEMIT after 80 editing steps. Besides, the hook layer causes some side effects in the locality of COUNTERFACT, but this circumstance is not found in the ZsRE dataset. It is worth noting that CoachHooK shows the most stable performance as the number of consecutive editing steps grows, showing the great potential of our method for consecutive editing.
Effect of the Balance Factor $\lambda$
Figure 5: Performance comparisons on initial five different values of $\lambda$.
Figure 6: Extension on the best three values of $\lambda$.
We test the effect of different $\lambda$, the balance factor between pre-training and newly updated associations. We first evaluate the CoachHooK with different $\lambda$ on 1k samples from ZsRE (Fig. 5). It seems that a small value of $\lambda=1,000$ would cause significant damage to all three metrics, especially the reliability and generality, since they experience a great drop as the consecutive steps increase. This may result from the overly high magnitude of the weight change caused by the low value of $\lambda$, which severely distorts the previously updated associations. Meanwhile, a too-high value of $\lambda=20,000$ also seems not to be a good choice, which gives rise to an overly small magnitude of the weight change so that it fails to deliver the new optimized values for keys. The cases of $\lambda=5000,10000,15000$ do not show an apparent difference, so we extend further the sample size to 10k (Fig. 6).
Extended results show that 5000 is not a good choice for large-consecutive editing steps, though it performs no worse than the other two in early 1k samples. The case of $\lambda=15,000$ ranks first in reliability and generality. Although it performs worse in locality compared to $\lambda=10,000$, the margin between them gradually narrows as the consecutive steps rise. Overall, we conclude that 15,000 would be a reasonable selection.
Effect of Editing Batch Size
Figure 7: Performance comparisons on different editing batch sizes.
The results show that while the three metrics decrease as the batch size rises, the margin could be negligible, denoting that our method possesses the mass-editing capacity.
More Analyses
Other detailed analyses including hyperparameters search, inference time and memory analysis are presented in Appendix C.
5 Related Work
Linear Associative Memory
Linear Associative Memory Kohonen (1972b); Anderson (1972b) treats the model weights $W$ as a set of associations between a set of keys and values. Many researchers have adopted this technique as a memory component in their methods. Schlag and Schmidhuber (2018); Shi et al. (2022); Li et al. (2023a); Chen et al. (2020); Schlag et al. (2021). Recent research Geva et al. (2021) finds that the transformer feed-forward layers are likely to be a key-value memory. This significantly helps researchers to understand what is happening inside the transformer black box and inspires researchers in the field of model editing to first locate where the knowledge is stored and then edit the corresponding parameters Meng et al. (2023, 2022).
Model Editing
Recent years have witnessed prosperous development in the field of model editing. According to Yao et al. (2023), the proposed methods so far can be generally classified into two groups, i.e., modify the model’s weight or not. The methods that do not directly alter the model weights generally follow two directions: they either employ an external memory or introduce additional adjustable parameters. Methods like T-Patcher Huang et al. (2023) and CaliNET Dong et al. (2022) apply new neurons that are responsible for specific mistakes in the last layer of the FFN model. GRACE Hartvigsen et al. (2022) introduces a timely adjusted code book to edit the model’s behavior. Another group of methods like SERAC Mitchell et al. (2022b) integrates an explicit external memory as edit descriptors to help editing scope recognition. Some other techniques like IKE Zheng et al. (2023), MeLLo Zhong et al. (2023), and MemPrompt Raffel et al. (2020) take advantage of the recent development of in-context learning and use carefully designed demonstration context prompts to edit the model. Cai et al. (2024) also discussed the possible transformation between context and parameter update. On the other hand, those directly altering the model’s weight either train a hyper-network to predict the change required by the edits or first locate corresponding parameters responsible for specific knowledge and then edit the located parameters. For example, Methods like KE De Cao et al. (2021), MEND Mitchell et al. (2022a), and MALMEN Tan et al. (2024) follow the idea of meta-learning and use the trained hypernetwork to predict the desired parameter change for editing. ROME Meng et al. (2022), MEMIT Meng et al. (2023), PMET Li et al. (2023b), KN Dai et al. (2022), and WilKE Hu et al. (2024) first locate the parameters responsible for knowledge and then using a designed optimization mechanism to shift the parameters.
Despite the good performance of the proposed methods, researchers investigated the harmful consequences of these methods. Yang et al. (2024) points out even few edits could result in the collapse of a model. Gupta et al. (2024) shows that scaling up the edits could lead to catastrophic forgetting of the previous edits. Gu et al. (2024) argues that editing could lead to the degradation of the model’s general capabilities. Although the rationality of these investigations is still under discussion, they reveal the potential risk that model editing techniques could bring to the base model.
6 Conclusion
This work introduces a novel model editing method, CoachHooK, which advocates the more practical consecutive batch model editing. CoachHooK uses an expanded editing mechanism to support consecutive editing and newly proposed hook layers to identify the editing scope. Compared to existing model editing methods, CoachHooK does not require large external memory nor extra training for meta-networks or classifiers. Instead, it adopts hook layers whose size remains fixed over time for storing associations. Comprehensive experiments are conducted to verify the method’s effectiveness over single-round and consecutive batch editing.
Limitations
Several aspects remain to be further investigated.
Other types of tasks
Notably, model editing techniques could be applied to various types of tasks. Specifically, besides factual knowledge editing, it can be applied to erase hallucinations, biases, privacy information, etc. However, the concentration of this paper is to explore the practicability of expanding the model editing application scenario to consecutive batch editing and investigate the potential bottleneck of corresponding methods under this scenario. Therefore, our experiment focuses on varying the scale of editing samples in factual knowledge editing tasks, as it is a relatively well-studied and universal evaluation task in model editing.
Model scale and architecture
Due to the limited computational resources, we cannot verify our method’s effectiveness in larger LLMs such as Llama-2 Touvron et al. (2023), and GPT-NEOX-20B Black et al. (2022). We focus on the decoder-only autoregressive models and do not include encoder-decoder structure models, as the autoregressive structures are the mainstream architecture nowadays OpenAI (2023); Touvron et al. (2023). Further, as stated by Yao et al. (2023), the weight matrix in some models like OPT-13B Zhang et al. (2022) is not invertible. However, such an issue can be relieved by adding a term $\beta I$ to the Eq. 14, where $\beta$ is a scalar expected to be small and $I$ is the identity matrix.
The shrink of $\alpha$
As more and more associations are integrated into the hook layer, the dynamically determined hyperparameter $\alpha$ will gradually shrink, meaning that an increasing number of vector entries in the original layer output will be swapped by the output from the hook layer, which is likely to lead the drop in locality. Nevertheless, such a problem can be alleviated by the newly designed updated mechanism (Eq. 13), which considers both previously updated and newly updated keys.
Appendix A Experiment Details
All baselines are implemented using the EasyEdit Wang et al. (2023) library.